OpenStudy (anonymous):

What is the simplified form of 3x+3/x-2 - x-4/x-2

OpenStudy (anonymous):

\[\frac{3x+3}{x-2} - \frac{x-4}{x-2} = \frac{3x+3-(x-4)}{x-2}\]

OpenStudy (anonymous):

thats how I'd start

OpenStudy (anonymous):

and then what would i do? im really lost